The energy of an electron in nth Bohr orbit of
hydrogen atom is

Grade:12

1 Answers

Arun
4 years ago

Energy of nth Bohr orbit for hydrogen like atoms can be calculated using the r/n -

En=-13.6 /n^2

So we get E1=-13.6

Lly E2=-3.4

Hence energy required for excitation of e- from first orbit to second orbit is

E2-E1=-13.6+3.4=-10.4